# LLM Agents for Education: Advances and Applications

### v1

@@ 3.1.2. Feedback Comment Generation (FCG)
Providing timely, relevant, and constructive feedback is a cornerstone of effective education. Teaching assistant agents can generate automated feedback comments on students’ assignments, quizzes, and projects. For example, Guo et al. (2024a) presents a training-free system for providing accurate feedback to students, which introduces a two-agent system. Specifically, Agent 1 generates initial feedback based on the students’ responses, while Agent 2 evaluates and refines this feedback to prevent over-praise and excessive inferences. Furthermore, Nair et al. (2024) designs a training strategy called PROF, which trains an automated LLM-based writing comment generator through reinforcement learning. This system adopts an iterative pipeline to simulate various student writing styles and incorporates a more advanced revision model (e.g., GPT-4) to provide the quality of the feedback as rewards. Similarly, SEFL (Zhang et al., 2025a) enhances feedback generation by having LLM agents role-play both students and teachers to generate data, which is then used to fine-tune models and improve feedback capabilities. These systems have also been deployed in real-world applications, such as FreeText (Matelsky et al., 2023), which pairs student responses with teacher-provided criteria, enabling the agent to identify strengths and weaknesses and provide targeted feedback for improvement. Beyond traditional feedback, advanced LLM agents are now capable of handling more complex, expertise-intensive tasks. For example, Du et al. (2024) explores the potential of LLM agents as assistants for natural language processing (NLP) paper reviewing tasks, while AAAR-1.0 (Lou et al., 2024) evaluates agents’ capabilities in areas such as equation inference, experiment design, paper weakness analysis, and review critique, revealing their potential in conducting sophisticated research tasks.
@@
3.1.3. Learning Resource Recommendation (LRR)
@@
To ensure that students access the most suitable learning resources aligned with their learning path and knowledge domain, it is crucial to develop an effective recommendation system. Zaiane (2002) first introduced recommendation systems into e-learning, utilizing web mining techniques to suggest online learning activities or shortcuts on course websites. Since then, the integration of LLM-based agents has opened new avenues for learning resource recommendations. These recommendations can be generated using both retrieval-based and generation-based methods. Retrieval-based methods involve agents accessing a database or their own memory to suggest existing resources—such as textbooks, research papers, or online content—based on student queries, past behavior, or content similarities (Shahzad et al., 2025; Li et al., 2024d). In contrast, generative methods create new learning content tailored to an individual student’s learning style, knowledge gaps, and interests (Moon et al., 2024). Moreover, to enhance the understanding and acceptance of recommended content, these agents need to provide reasons for their recommendations. Explaining the rationale behind suggestions fosters trust and enables students to make more informed decisions about the resources they engage with. For example, Abu-Rasheed et al. (2024a, b) incorporate knowledge graphs to recommend human-curated sources of information. This approach not only enhances the interpretability of recommendations but also reduces the risk of generating misinformation, thereby improving the quality of the learning experience.
@@
Future work should focus on integrating agents with adaptive learning systems to dynamically adjust recommendations based on real-time student performance. Also, a hybrid approach combining generative and retrieval-based methods could enhance both accuracy and diversity. Furthermore, incorporating multi-modal content—including interactive media, video, and immersive simulations—could further enrich the learning experience.
3.2. Agent for Student Support
@@
LLM agent-based student support systems enable independent, personalized learning by providing real-time academic assistance without requiring direct teacher involvement. Unlike traditional rule-based learning systems, they deliver interactive, tailored feedback, allowing students to learn at their own pace. Broadly, their core functionalities include adaptive learning3.2.1), knowledge tracing3.2.2), and error correction and detection3.2.3).
3.2.1. Adaptive Learning (AL)
LLM agents can automate tasks, making it possible to build self-sustaining adaptive learning systems without teacher involvement. These systems operate within an action-memory framework, continuously refining instruction based on student profiling, content selection, and difficulty adaptation. The agent maintains a structured student profile, dynamically selecting materials based on past interactions. After each task, it analyzes student behavior and adjusts teaching strategies accordingly, enhancing personalization and responsiveness.
@@ 3.2.1. Adaptive Learning (AL)
Several implementations exemplify this adaptive approach. EduAgent (Xu et al., 2024a) introduces a structured profiling mechanism comprising four distinct cognitive patterns: gaze behavior linked to physiological memory, motor behavior mapped to motor memory, cognitive state associated with cognitive memory, and post-course assessments contributing to knowledge memory. This structured representation enhances adaptive decision-making by providing a multi-faceted view of student learning states. Chen et al. (2024b) propose a system consisting of interaction, reflection, and reaction, with each component composed of specific LLM tools and memory modules. Furthermore, a meta-agent is introduced to control the information flow through these agents.
@@
Building on these adaptive mechanisms, LLM agents operate within a state-action framework, allowing them to dynamically adjust instructional strategies in response to a student’s evolving cognitive and affective state. The state space in these systems encodes the student’s cognitive and affective profile, while the action space governs how the agent adjusts its teaching approach. A typical cognitive state representation includes tracking a student’s knowledge proficiency, comprehension levels, and misconceptions (Park et al., 2024; Liu et al., 2024c, a), allowing the agent to tailor explanations, adjust difficulty levels, and reinforce concepts dynamically. Recent studies highlight the importance of affective state modeling, as emotional factors such as motivation, interest, and self-efficacy significantly influence learning outcomes. For instance, Park et al. (2024) propose an affective state model that enables agents to adjust feedback tone, provide encouragement, and regulate pacing to maintain engagement. Another crucial dimension of adaptation involves learning preferences and personality traits. These studies integrate personality with memory design, tracking the psychological state of the students. Wang et al. (2025b) integrate learning preferences into state modeling, recognizing that students process information differently depending on instructional format and modality. Adapting content to these preferences enhances retention and learning efficiency. Moreover, Liu et al. (2024c) apply the Big Five personality model (Roccas et al., 2002) to personalize tutoring strategies, acknowledging that individual differences shape learning experiences (Sonlu et al., 2024).
@@
Recent approaches explore multi-agent systems for adaptive learning, where specialized agents collaborate to enhance personalization. For example, Wang et al. (2025b) design five agents—Gap Identifier, Learner Profiler, Dynamic Learner Simulator, Learning Path Scheduler, and Content Creator—to deliver goal-oriented, personalized instruction. Similarly, OATutor (Pardos et al., 2023) provides an experimental platform for modular adaptive learning, allowing researchers to design scalable, domain-general tutoring agents.
3.2.2. Knowledge Tracing (KT)
Knowledge tracing is a fundamental component of intelligent tutoring systems, enabling the tracking of a learner’s evolving understanding over time to predict future performance and knowledge retention. Traditional knowledge tracing methods often rely on statistical models or deep learning approaches to estimate student mastery levels. In contrast, LLM agents introduce a more dynamic and personalized approach, leveraging their ability to process natural language interactions, infer conceptual mastery, and adapt instructional strategies accordingly.
@@ 3.2.2. Knowledge Tracing (KT)
Recent advancements have explored multi-agent frameworks for knowledge tracing. For instance, Yang et al. (Yang et al., 2024a) propose a multi-agent system with three specialized agent roles: administrator, judger, and critic. In this framework, the administrator delegates knowledge tracing tasks to judgers, who collaborate through discussions to assess the student’s cognitive state. The critic agent then evaluates the outcome and determines whether the assessment criteria are met, ensuring a structured yet flexible knowledge tracing process. Other agent-based approaches explore alternative strategies for modeling student knowledge. Xu et al. (2024a) propose simulating students as different personas, allowing agents to adaptively trace knowledge progression based on varied learning profiles. Meanwhile, Scarlatos et al. (2025) employ dialogue-driven interactions to probe students’ conceptual boundaries, using conversational exchanges to refine knowledge estimation dynamically.
3.2.3. Error Correction and Detection (ECD)
@@
Error detection and correction are critical components of intelligent tutoring systems, enabling students to refine their understanding by receiving real-time, context-aware feedback. LLM agents can identify errors in student responses across various domains, including academic writing, programming, and mathematical reasoning. By dynamically adapting feedback to the learner’s proficiency level, these agents serve as intelligent reviewers, debuggers, and writing assistants (Ye et al., 2022; Li et al., 2024c).
@@
Agent-based systems leverage state representations and adaptive inference mechanisms to track error patterns and misconceptions dynamically (Park et al., 2024; Liu et al., 2024c; Wang et al., 2025b; Bouzenia et al., 2024). Recent advancements extend this capability into the multi-modal domain, incorporating direct analysis of student-generated drafts. Xu et al. (2024b) propose a multimodal LLM framework that processes handwritten or digitally drafted student work. The system first extracts and converts draft content into natural language, enabling the LLM-based agent to interpret and analyze handwritten responses. The agent then provides indirect yet effective instructional feedback, guiding students toward self-correction and deeper comprehension. Moreover, CoT Rerailer (Wan et al., 2024) designs a derailment identification process and a rerailment process to conduct error detection. Zhang et al. (2025b) propose the MathCCS benchmark and develop a sequence error analysis framework, adopting multi-agent collaboration. As the first benchmark for multimodal error detection, ErrorRadar (Yan et al., 2024c) can provide a data foundation for multimodal agents in error detection task.
4. Domain-Specific Educational Agents
Recent research on LLM agents in education has also shown growing interest in domain-specific applications. We explore their use in science learning, language learning, and professional development, focusing on their algorithmic frameworks, agentic designs, and relevant datasets and benchmarks.
4.1. Agent for Science Learning
An agent for science learning is an intelligent system powered by LLMs, designed to assist students in acquiring and applying scientific knowledge through personalized, interactive experiences (Yan et al., 2025b; Raihan et al., 2025; Ng et al., 2024; Brown and Cruz Castro, 2025). The significance of these agents in education lies in their ability to offer tailored feedback, enhance conceptual understanding, and promote active engagement with complex scientific ideas. In the following sections, we explore the impact of LLM agents in four key scientific disciplines: mathematics4.1.1), physics4.1.2), chemistry4.1.3), and biology4.1.4), as well as their broader contributions to general scientific discovery4.1.5).
4.1.1. Mathematics
In mathematics, LLM agents provide substantial support by helping students navigate complex problems and reinforcing their understanding of abstract concepts (Yan et al., 2024b; Xiong et al., 2024; Swan et al., 2023; Yan et al., 2024c; Wu, 2025; Mitra et al., 2024). For instance, Gou et al. (2023) introduce TORA (Tool-integrated Reasoning Agents), a framework that integrates natural language reasoning and program-based tool use to handle mathematical reasoning. MathAgent (Yan et al., 2025a) similarly proposes Mixture-of-Math-Agent framework to address multimodal error detection in real-world K-12 scenarios, and flexibly transform the visual information of different types of questions into forms that are more easily understood by LLMs (e.g., converting plane geometry images into formalized expression). Additionally, MathChat (Wu et al., 2023) serves as a conversational mathematical problem-solving agent, which consists of a chat-based LLM agent and a tool-based user agent. Furthermore, Xiong et al. (2024) propose to use reinforcement learning from human feedback (RLHF) to further improve tool-integrated agents for mathematical problem-solving, and formulate this method as a Markov decision process, distinguishing it from the typical contextual bandit approach used in RLHF. Besides, MACM (Lei et al., 2025) discuss the limitations of LLMs in handling complex mathematical logical deduction, thus introducing a multi-agent system, which comprises three interactive agents: Thinker, Judge, and Executor.
4.1.2. Physics
In the field of physics, LLM agents help students make sense of challenging concepts and offer interactive tools to simulate physical phenomena (Pang et al., 2024; Mower and Bou-Ammar, 2025; Barman et al., 2025; Feng et al., [n. d.]; Jiang and Jiang, 2024; Yan and Lee, 2024). Wang et al. (2023) introduce NEWTON, the first pipeline and benchmark to explore the physical reasoning abilities of LLMs. Furthermore, Kortemeyer (2023) describe a case study exploring if an LLM agent can pass an introductory calculus-based physics course. In addition, Physics Reasoner (Pang et al., 2024), a novel knowledge-augmented framework for physics problem-solving, leverages a comprehensive formula set and detailed checklists to ensure accuracy and completeness. It can serve as an agent consisting of three stages - problem analysis, formula retrieval, and guided reasoning. Besides, Ma et al. (2024c) describe the Scientific Generative Agent (SGA), a bilevel optimization framework designed for physical scientific discovery, and highlight the use of LLMs for generating and revising scientific hypotheses and implementing an exploit-and-explore strategy.
4.1.3. Chemistry
Chemistry education also benefits greatly from LLM agents, which can explain molecular structures, chemical reactions, and experimental processes in an engaging and interactive way (Ramos et al., 2025; M. Bran et al., 2024; Yu et al., 2024; Guo et al., 2023; Tsai et al., 2023). For example, ChemCrow (M. Bran et al., 2024) is the first LLM chemistry agent capable of autonomous planning and execution of chemical syntheses, including an insect repellent and three organocatalysts. Yu et al. (2024) further present ChemAgent, an enhanced chemistry agent improved over ChemCrow, with a focus on two essential cognitive abilities of chemistry problem-solving: reasoning and grounding. Besides, Curie (Kon et al., 2025) is an agent framework aimed at incorporating rigor into the experimentation process via three core elements: an intra-agent rigor module to boost reliability, an inter-agent rigor module to ensure systematic control, and an experiment knowledge module to improve interoperability. Recent studies have explored the capabilities of LLMs in complex chemical discovery (Yang et al., 2024c; Ruan et al., 2024; Sprueill et al., 2024; Moret et al., 2023; Jablonka et al., 2024), and their potential can be advanced by leveraging the interactivity of agent-based tool use and the flexibility of planning strategies (Song et al., 2024; Ramos et al., 2025).
4.1.4. Biology
In biology, LLM agents enhance learning by offering detailed explanations of biological processes and providing interactive experiences to explore living systems (Yan et al., 2025b; Bhattacharya et al., 2024; Sripathi et al., 2024; Zhao et al., 2025; Gao et al., 2024). For example, ProtChat (Huang et al., 2024b) is a multi-agent tool leveraging GPT-4 and Protein Language Models for seamless protein analysis automation, thus evolutionizing the complexities of protein sequence interpretation. ProtAgents (Ghafarollahi and Buehler, 2024a) is introduced as a multi-agent modeling framework that combines state-of-the-art LLMs with diverse tools to tackle protein design and analysis. It consists of a team of agents: User, Planner, Assistant, Critic, and Group Chat Manager. Besides, Shen et al. (2024) present TourSynbio-Agent, an innovative agent framework that leverages TourSynbio-7B’s protein understanding ability to perform various protein engineering tasks, such as mutation analysis, inverse folding, and visualization.
4.1.5. General Scientific Discovery
LLM agents support general scientific discovery by assisting students in data interpretation, hypothesis testing, and creative problem-solving (Yan et al., 2025b; Chen et al., 2024a; Ghafarollahi and Buehler, 2024b; Chen et al., 2025b; Schmidgall et al., 2025). These LLM agents, such as PaperQA (Lála et al., 2023), PaperQA2 (Skarlinski et al., 2024), OpenScholar (Asai et al., 2024), SciAgents (Ghafarollahi and Buehler, 2024b), TheoremExplainAgent (Ku et al., 2025), and LLM-SR (Shojaee et al., 2024), can analyze complex scientific datasets, helping students uncover patterns and trends that may not be immediately apparent. In addition, Narayanan et al. (2024) present Aviary, an extensible gymnasium for language agents for three challenging scientific tasks: manipulating DNA constructs for molecular cloning, answering research questions by accessing scientific literature, and engineering protein stability. Furthermore, both SciAgent (Ma et al., 2024a) and SciToolAgent (Chen et al., 2025b) extend to a tool-augmented scientific reasoning setting with the help of domain-specific tools. Besides, Agent Laboratory (Schmidgall et al., 2025) emerges as an agent framework that automates the research process of three phases (Literature Review, Experimentation, and Report Writing) via various LLM agents (PhD, Postdoc, ML Engineer, etc.).
4.2. Agent for Language Learning
The integration of LLM agents into language learning is revolutionizing how core competencies—reading, writing, listening, and speaking—are taught and practiced (Ye et al., 2025b). These skills form the foundation of effective communication and language acquisition, and recent advancements in LLM-based agents have significantly enhanced how learners interact with and acquire these skills. Below, we introduce recent advancements in each subdomain, highlighting the role of LLM agents in enhancing pedagogical outcomes in language acquisition for students and second language (L2) speakers through engaging and adaptive approaches (Huang et al., 2023b; Ye et al., 2023a, b).
4.2.1. Reading
Reading comprehension is a vital component of language learning, and LLM agents are playing an increasingly important role in enhancing students’ reading abilities. For instance, ExpertEase (Mo and Hu, 2024) employs a multi-agent framework to adapt documents for grade-specific audiences, simulating expert-teacher-student collaboration to enhance comprehension. AgentSimp (Fang et al., 2025) tackles document-level simplification by leveraging multiple agents with distinct roles to ensure coherence and accessibility. Additionally, various LLMs (Adetayo et al., 2024) have been used as academic reading companions, demonstrating improved engagement and understanding of complex qualitative texts in educational settings (Chen and Leitch, 2024).
4.2.2. Writing
The development of writing skills has benefited significantly from NLP tasks like explainable grammatical error correction (EXGEC) (Ye et al., 2024b, 2025a; Zou et al., 2025) and automatic essay scoring (AES) systems (Su et al., 2025). Weaver (Wang et al., 2024b), a family of LLMs fine-tuned for writing tasks (Ye et al., 2023c, 2024c), outperforms generalist LLMs like GPT-4 in generating human-like narratives. Moreover, Weaver natively supports retrieval-augmented generation (RAG) and function calling, serving as a qualified foundational model for LLM agents. For interactive feedback on student essays, CAELF (Hong et al., 2024) introduces a multi-agent framework that enables interactive essay feedback. By combining Teaching-Assistant agents’ evaluations with teacher-agent arbitration, students can contest grades and engage with the feedback, addressing the “black box” limitations of traditional automated scoring. Inspired by the process of human debate, Debate-to-Write (Hu et al., 2025a) construct a persona-based multi-agent framework that can enable agents to collaboratively debate, discuss ideas, and form a comprehensive plan for argument writing.
4.2.3. Translation
LLM agents demonstrate remarkable advancements in both simultaneous (Koshkin et al., 2024; Guo et al., 2024c) and literary translation (Cheng et al., 2024; Škobo and Petričević, 2023). Translation tasks benefit from LLM agents through their ability to integrate specialized tools and orchestrate multi-agent collaboration. Agent-SiMT (Guo et al., 2024b) combines the decision-making capabilities of a Simultaneous Machine Translation (SiMT) policy agent with the generative power of a translation agent, achieving state-of-the-art performance in simultaneous translation by dynamically balancing reading and generation actions. For literary translation, TransAgents (Wu et al., 2024) employs a multi-agent framework to replicate the complex workflows of human translation teams, addressing cultural nuances and stylistic challenges through collaborative reasoning. This approach not only improves translation quality but also extends LLM applications to linguistically and culturally rich domains. These contributions underscore the importance of tool use and agent collaboration in advancing translation education (Zheng et al., 2024).
4.2.4. Storytelling
Storytelling applications leverage LLM agents to create immersive and interactive learning experiences (Simon and Muise, 2022). STARie (Li and Xu, 2023), a peer-like embodied conversational agent, integrates multimodal tools such as speech synthesis and facial animation to scaffold children’s storytelling, fostering narrative creativity and oral communication skills (Beredo and Ong, 2021; Cassell, 2022). StoryAgent (Sohn et al., 2024) combines top-down story drafting with bottom-up asset generation to transform simple prompts into coherent, multi-modal digital narratives. By automating complex storytelling workflows (Liem et al., 2023), it democratizes content creation and enhances engagement in language learning. LLaMS (Zhang et al., 2025c), a multi-modal agent framework, is designed to generate multi-modal human-level stories characterized by expressiveness and consistency, incorporating the Story-Adapter module for long image sequence illustration. These systems demonstrate the potential of LLM agents to support both cognitive and creative aspects of language education for children.
4.2.5. Speaking
LLM agents are revolutionizing spoken language education by integrating reasoning and multi-agent collaboration to build adaptive dialogue systems (Liu et al., 2024b; Balan et al., 2024). ELLMA-T (Pan et al., 2024) employs contextual reasoning and role-playing in social VR environments to provide personalized feedback and language assessments, enabling learners to practice speaking in realistic scenarios (Lim et al., 2024; Li et al., 2025b). SpokenWOZ (Si et al., 2023) introduces a large-scale benchmark for task-oriented spoken dialogue, highlighting the importance of reasoning and multi-turn interaction in addressing real-world conversational challenges. FurChat (Cherakara et al., 2023), an embodied conversational agent, combines verbal and non-verbal communication cues to simulate natural interactions, making it a valuable tool for improving speaking skills through immersive and realistic practice. By employing multimodal signals such as speech and gestures, SpeechAgents (Zhang et al., 2024c) enhances the authenticity of dialogue simulations, capturing consistent content, natural rhythm, and rich emotional expression. Through Multi-Agent Tuning (Liang et al., 2024), it optimizes LLM capabilities for large-scale simulations involving up to 25 agents, enabling applications like drama creation and audio novel generation.
4.3. Agent for Professional Development
Agents for professional development harness the capabilities of LLMs to offer scalable, adaptive, and context-aware learning experiences tailored to domain-specific needs. This section summarizes how recent studies develop agents to revolutionize professional training in fields including medical4.3.1), computer science4.3.2), and law education4.3.3).
4.3.1. Medical Education
The deployment of LLM agents in healthcare has created new opportunities for personalized, interactive, and scalable systems, with several health agents introduced (Shusterman et al., 2025) such as Talk2Care (Yang et al., 2024d) and openCHA (Abbasian et al., 2023). Additionally, Abd-Alrazaq et al. (2023) highlight the educational potentials of LLMs in crafting personalized curricula, adaptive learning plans, and dynamic assessment tools for medical education, while concurrently addressing challenges including algorithmic bias, misinformation, and privacy issues. MEDCO (Wei et al., 2024), a multi-agent system, has the capacity to replicate real-world medical training environments through agent collaboration with virtual patients, expert physicians, and radiologists, enhancing interdisciplinary learning and peer interaction. Furthermore, Abbasian et al. (2023) introduce openCHA as a personalized LLM-powered framework that integrates external resources and orchestrates multi-step problem-solving for complex healthcare queries (Ye et al., 2024a), emphasizing tool use and action planning. Beyond traditional education, Steenstra et al. (2024) explore LLMs in creating fantasy narrative games for adolescent health education, demonstrating the agents’ ability to generate engaging, doctor-validated content that enhances knowledge retention through gamification. Li et al. (2024b) present Agent Hospital, a simulation environment where LLM-driven agents evolve through autonomous interactions, demonstrating significant improvements in medical reasoning and performance on benchmarks like MedQA (Jin et al., 2021) after treating thousands of simulated patients. Collectively, these investigations highlight the versatility of LLM agents within medical education, demonstrating their abilities in reasoning, collaboration, tool integration, and adaptive learning to effectively address a broad spectrum of educational and clinical challenges (Karabacak and Margetis, 2023; Tian et al., 2024b; Ullah et al., 2024).
4.3.2. Computer Science Education
An agent for computer science (CS) education greatly enhances learning by providing personalized guidance on coding, debugging, and understanding CS principles (Ma et al., 2024b; Lee and Song, 2024; Kosar et al., 2024; Liu et al., 2024d). For example, CodeAgent (Zhang et al., 2024b) serves as an LLM agent framework for repo-level code generation, incorporating external tools such as WebSearch and DocSearch. Recent studies have demonstrated the potential of agent-based code generation systems such as ToolCoder (Ding et al., 2025), SWE-agent (Yang et al., 2025), AgentCoder (Huang et al., 2023c), and MapCoder (Islam et al., 2024), which can significantly enhance students’ coding efficiency (Jin et al., 2024a; Wang et al., 2025a; Frankford et al., 2024). Furthermore, Bassner et al. (2024) introduce Iris, an LLM-driven virtual tutor designed to offer personalized, context-aware assistance to CS students within the interactive learning platform Artemis. Besides, Chen et al. (2024d) propose Learning-by-Teaching (LBT) as an effective pedagogical strategy for CS education, and leverage the advantages of LLM agents (e.g., contextual conversation & learning from demonstrations).
4.3.3. Law Education
LLM agents leverage pre-trained legal knowledge, interactive capabilities, and reasoning skills to support law education through judicial interpretation, moot court simulation, and case analysis (Chen et al., 2024c; Nelson, 2024; Lai et al., 2024; Yuan et al., 2024). However, evaluations from LawBench (Fei et al., 2023) and LegalBench (Guha et al., 2023) reveal that LLMs struggle with legal knowledge application and judicial aid. LegalAgentBench (Li et al., 2024a) further highlights their limitations in multi-hop reasoning and defense statement writing, showing that LLM agents require significant improvements to effectively assist in complex legal tasks. Despite these challenges, LLM agents are emerging as valuable tools for moot court simulations, a crucial component of legal reasoning and advocacy training. DeliLaw (Xie et al., 2024a) enhances law education by integrating legal and case retrieval modules, enabling students to practice legal research, case analysis, statutory interpretation, and mock consultations. LawLuo (Sun et al., 2024) applies a multi-agent framework with retrieval-augmented generation to simulate multi-turn legal consultations, improving personalization and ambiguity handling. Similarly, AgentCourt (Chen et al., 2024c) and AgentsCourt (He et al., 2024a) simulate courtroom interactions and judicial decision-making, providing a realistic training ground for law students. AgentsBench (Jiang and Yang, 2024) extends this by offering multi-agent legal reasoning and case analysis, further advancing AI-driven legal education.
5. Challenges & Future Directions

### v2

@@ 3.2 Feedback Comment Generation
For example, Guo et al. (2024c) leverage multi-agent communication to provide accurate feedback to students through a two-agent system. Specifically, Agent 1 generates initial feedback based on the students’ responses, while Agent 2 evaluates and refines this feedback to prevent over-praise and excessive inferences. Furthermore, Nair et al. (2024) design a training strategy called PROF, which trains an automated LLM-based writing comment generator through reinforcement learning. This system adopts an iterative pipeline to simulate various student writing styles and incorporates a more advanced revision model (e.g., GPT-4) to provide the quality of the feedback as rewards. Similarly, SEFL (Zhang et al., 2025a) enhances feedback generation by having LLM agents role-play both students and teachers to generate data, which is then used to fine-tune models and improve feedback capabilities. These systems have also been deployed in real-world applications, such as FreeText (Matelsky et al., 2023), which pairs student responses with teacher-provided criteria, enabling the agent to identify strengths and weaknesses and provide targeted feedback for improvement. Beyond traditional feedback, advanced LLM agents are now capable of handling more complex, expertise-intensive tasks. For example, Du et al. (2024) explores the potential of LLM agents as assistants for natural language processing paper reviewing tasks, while AAAR-1.0 (Lou et al., 2024) evaluates agents’ capabilities in areas such as equation inference, experiment design, paper weakness analysis, and review critique, revealing their potential in conducting research tasks.
@@
3.3 Curriculum Design
To ensure that students follow personalized learning paths aligned with their knowledge level and domain, it is crucial to develop effective curriculum design strategies. This is a complex task that relies on multiple LLM agent capabilities, including memory, tool use, planning, personalization, and explainability.
@@ 3.3 Curriculum Design
Zaiane (2002) first introduces recommendation systems into e-learning, utilizing web mining techniques to suggest online learning activities or shortcuts on course websites. The integration of LLM-based agents has recently enabled more sophisticated curriculum design through dynamic sequencing and content adaptation. Curricula can be built using either retrieval-based or generation-based methods. Retrieval-based methods involve agents accessing a database or their own memory to suggest existing resources—such as textbooks, research papers, or online content—based on student queries, past behavior, or content similarities (Shahzad et al., 2025; Li et al., 2024c). In contrast, generative methods create new learning content tailored to an individual student’s learning style, knowledge gaps, and interests (Moon et al., 2024). Moreover, to enhance the understanding and acceptance of recommended content, these agents need to provide reasons for their recommendations. Explaining the rationale behind suggestions fosters trust and enables students to make more informed decisions about the resources they engage with. For example, Abu-Rasheed et al. (2024a, b) incorporate knowledge graphs to guide curriculum pathways toward curated and trustworthy sources. This approach not only enhances the interpretability of recommendations but also reduces the risk of generating misinformation, thereby improving the quality of the learning experience.
@@
Looking ahead, future work should focus on integrating agents with adaptive learning systems to dynamically adjust recommendations in response to real-time student performance. In addition, a hybrid approach that combines generative and retrieval-based methods may improve both the accuracy and diversity of curriculum content. Also, incorporating multimodal resources such as interactive media, video, and immersive simulations could further enhance the learning experience.
4 Agent for Student Support
@@
LLM agent-based student support systems aim to provide real-time personalized assistance without requiring direct teacher involvement. Unlike traditional rule-based systems, these agents offer interactive and adaptive feedback, enabling students to progress at their own pace. Core educational tasks in student support include adaptive learning (§4.1), knowledge tracing (§4.2), and error correction and detection (§4.3).
4.1 Adaptive Learning
LLM agents offer the potential to build self-sustaining adaptive learning systems that operate without direct teacher involvement. These systems dynamically tailor instruction based on student performance, enabling personalized learning at scale. Building on the abilities of memory and personalization, agents are able to maintain and update a structured representation of the learner, commonly referred to as a student profile. This profile informs content selection, pacing, and feedback strategies.
@@ 4.1 Adaptive Learning
Several implementations exemplify this adaptive approach. GenAL Lv et al. (2025) integrates external tools such as automated programs and web searches to construct comprehensive student profiles and inform instructional planning. Based on these profiles, the agent assigns tasks that align with the student’s current knowledge level and updates its memory dynamically. EduAgent Xu et al. (2024a) introduces a structured profiling mechanism comprising four distinct cognitive patterns: gaze behavior linked to physiological memory, motor behavior mapped to motor memory, cognitive state associated with cognitive memory, and post-course assessments contributing to knowledge memory. This structured representation enhances adaptive decision-making by providing a multi-faceted view of student learning states. Chen et al. (2024c) propose a system consisting of interaction, reflection, and reaction, with each component composed of specific LLM tools and memory modules. Furthermore, a meta-agent is introduced to control the information flow through these agents.
@@
Some recent research focuses on modeling students through modular memory or state components that capture cognitive, affective, and psychological dimensions. These representations are often formalized within a state–action framework, where the state space encodes learner traits and the action space governs instructional adaptations. A typical cognitive state representation includes tracking a student’s knowledge proficiency, comprehension levels, and misconceptions (Park et al., 2024; Liu et al., 2024c, a), allowing the agent to tailor explanations, adjust difficulty levels, and reinforce concepts dynamically. Recent studies highlight the importance of affective state modeling, as emotional factors such as motivation, interest, and self-efficacy significantly influence learning outcomes. For instance, Park et al. (2024) propose an affective state model that enables agents to adjust feedback tone, provide encouragement, and regulate pacing to maintain engagement. Another crucial dimension of adaptation involves learning preferences and personality traits. These studies integrate personality with memory design, tracking the psychological state of the students. Wang et al. (2025b) integrate learning preferences into state modeling, recognizing that students process information differently depending on instructional format and modality. Adapting content to these preferences enhances retention and learning efficiency. Moreover, Liu et al. (2024c) apply the Big Five personality model (Roccas et al., 2002) to personalize tutoring strategies, acknowledging that individual differences shape learning experiences (Sonlu et al., 2024).
@@
Emerging works also explore multi-agent systems for adaptive learning, where specialized agents collaborate to enhance personalization. For example, Wang et al. (2025b) design five agents—Gap Identifier, Learner Profiler, Dynamic Learner Simulator, Learning Path Scheduler, and Content Creator—to deliver goal-oriented, personalized instruction. Similarly, OATutor Pardos et al. (2023) provides an experimental platform for modular adaptive learning, allowing researchers to design scalable, domain-general tutoring agents.
4.2 Knowledge Tracing
Knowledge tracing is essential for monitoring a learner’s evolving understanding and predicting future performance. While traditional methods use statistical or deep learning models to estimate mastery, LLM agents offer a more dynamic, personalized approach by leveraging natural language understanding and adaptive instruction.
@@ 4.2 Knowledge Tracing
Recent advancements have explored multi-agent frameworks for knowledge tracing. For instance, Yang et al. Yang et al. (2024b) propose a multi-agent system with three specialized agent roles: administrator, judger, and critic. In this framework, the administrator delegates knowledge tracing tasks to judgers, who collaborate through discussions to assess the student’s cognitive state. The critic agent then evaluates the outcome and determines whether the assessment criteria are met, ensuring a structured yet flexible knowledge tracing process. Other agent-based approaches explore alternative strategies for modeling student knowledge. Xu et al. (2024a) propose simulating students as different personas, allowing agents to adaptively trace knowledge progression based on varied learning profiles. Meanwhile, Scarlatos et al. (2025) employ dialogue-driven interactions to probe students’ conceptual boundaries, using conversational exchanges to refine knowledge estimation dynamically.
4.3 Error Correction and Detection
@@
Error detection and correction help students refine their understanding through real-time, context-aware feedback. LLM agents can identify mistakes across domains such as writing, programming, and math, and adapt feedback to the learner’s proficiency, acting as intelligent reviewers and assistants Ye et al. (2022); Li et al. (2024d).
@@
Agent-based systems leverage state representations and adaptive inference mechanisms to track error patterns and misconceptions dynamically (Park et al., 2024; Liu et al., 2024c; Wang et al., 2025b; Bouzenia et al., 2024). Recent advancements extend this capability into the multi-modal domain, incorporating direct analysis of student-generated drafts. Xu et al. (2024b) propose a multimodal LLM framework that processes handwritten or digitally drafted student work. The system first extracts and converts draft content into natural language, enabling the agent to interpret and analyze handwritten responses. The agent then provides indirect yet effective instructional feedback, guiding students toward self-correction and deeper comprehension. Moreover, CoT Rerailer Wan et al. (2024) designs a derailment identification process and a rerailment process to conduct error detection when solving math questions. Zhang et al. (2025d) propose the MathCCS benchmark and introduce a sequence error analysis framework that leverages multi-agent collaboration. As the first benchmark for multimodal error detection, ErrorRadar Yan et al. (2024c) provides a valuable data foundation for developing multimodal agents in this task.
5 Challenges and Future Directions